OpenStudy (anonymous):

HELP: You are shopping for jeans. City Express sells 3 pairs of jeans for 61$. Denim Planet sells 2 pairs of jeans for $73. New threads sells 4 pairs of jeans for $110. Which has the best

OpenStudy (anonymous):

best what?

OpenStudy (anonymous):

61/3 = about 20 73/2 = about 36 110/4 = about 27

OpenStudy (anonymous):

Compare the price per pair of jeans

OpenStudy (anonymous):

It's the first one, City express

OpenStudy (anonymous):

The best deal would be at City Express. There, a pair of jeans cost $20, while at Denim Planet a pair costs$36.50 and at New Thread a pair costs $27.50

OpenStudy (anonymous):

City express sells each jean for about $20 (61/3) Denim planet sells theirs for about $37 (73/2) New threads sells theirs for about $27 (110/4) City express is the best